OpenStudy (anonymous):

Please help! Find the eccentricity of this equation: (x-5)^2=4(x-5)

OpenStudy (anonymous):

\[(y-5)^{2}=4(x-5)\]

ganeshie8 (ganeshie8):

start by figuring out the type of conic section the equation represents

OpenStudy (anonymous):

i know that\[e=\frac{ c }{ a }\] is it a parabola? because only one variable is squared when distributing?

ganeshie8 (ganeshie8):

Yes ! it is a parabola, and what do you know about eccentricity of parabola ?

ganeshie8 (ganeshie8):

http://www.mathsisfun.com/geometry/eccentricity.html

OpenStudy (anonymous):

y=1 if its a parabola

ganeshie8 (ganeshie8):

thats right !!

OpenStudy (anonymous):

so i set \[\frac{ c }{ a }\] equal to 1?

ganeshie8 (ganeshie8):

eccentricity of parabola is 1. we're done, right ?

OpenStudy (anonymous):

what about c/a?

ganeshie8 (ganeshie8):

I think you can use that formula for ellipses/hyperbolas

OpenStudy (anonymous):

\[(y-5)^{2}=4(x-5)\] \[y ^{2}+25=4x-20\] how do i put it into one of those formulas?

OpenStudy (anonymous):

oh so like (y+5)^2-4(x+20)=1 kind of thing?

OpenStudy (anonymous):

\[(y+5)^{2}-4(x+20)=1\]

OpenStudy (anonymous):

that gives me the center...can i find a from the center?

OpenStudy (anonymous):

dont i need vertices to find a and b?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

eccentricity = \(\large \dfrac{\text{distance between a point to Focus}}{\text{distance between a point to Directrix}}\)

ganeshie8 (ganeshie8):